Alden works as an astronaut who has been sent to the International Space Station (ISS). At one instance while he is at rest, Alden throws a 5.0-kg bowling ball away from him at a rate of 2.0 m/s relative to ISS. At what speed and direction will he start to move after he releases the ball? *


1
Expert's answer
2021-01-23T10:28:50-0500

According to the law of conservation of momentum


"Mv+mu=0\to u=-\frac{Mv}{m}"


If the mass of Aldene "76(kg)", we get


"u=-\frac{Mv}{m}=-\frac{5\cdot 2}{76}=-0.13(m/s)" opposite to the direction of movement of the ball. Answer